### v2

# The Why Behind the Action: Unveiling Internal Drivers via Agentic Attribution

Abstract
Large Language Model (LLM)-based agents are widely used in real-world applications such as customer service, web navigation, and software engineering. As these systems become more autonomous and are deployed at scale, understanding why an agent takes a particular action becomes increasingly important for accountability and governance. However, existing research predominantly focuses on failure attribution to localize explicit errors in unsuccessful trajectories, which is insufficient for explaining the reason behind agent behaviors. To bridge this gap, we propose a novel framework for general agentic attribution, designed to identify the internal factors driving agent actions regardless of the task outcome. Our framework operates hierarchically to manage the complexity of agent interactions. Specifically, at the component level, we employ temporal likelihood dynamics to identify critical interaction steps; then at the sentence level, we refine this localization using perturbation-based analysis to isolate the specific textual evidence. We validate our framework across a diverse suite of agentic scenarios, including standard tool use and subtle reliability risks like memory-induced bias. Experimental results demonstrate that the proposed framework reliably pinpoints pivotal historical events and sentences behind the agent behavior, offering a critical step toward safer and more accountable agentic systems. Codes are available at https://github.com/AI45Lab/AgentDoG.
Machine Learning, ICML
1 Introduction
Large Language Model (LLM)-based agents have emerged as a powerful paradigm for automated task execution, demonstrating impressive capabilities in various domains (Yao et al., 2022; Qian et al., 2023; Guo et al., 2024; Wang et al., 2024a). By integrating long-term memory, external tools, and iterative reasoning loops, these agents are widely applied to dynamic environments and complex tasks ranging from software engineering to web navigation and scientific discovery (Jin et al., 2024; Ning et al., 2025; Ren et al., 2025; Guo et al., 2025). As agents become more capable and autonomous, failures in agentic systems have attracted increasing attention. Recent advances in Agentic Failure Attribution focus on identifying where an agent fails during task execution (Zhang et al., 2025c, b; Zhu et al., 2025). These approaches typically analyze completed trajectories after an explicit failure signal, such as an incorrect answer, an execution error, or task non-completion. By localizing faulty steps, they provide effective tools for debugging and improving system reliability.
Figure 1: Paradigm shift from failure attribution to agentic attribution. Top: Traditional failure attribution targets unsuccessful trajectories, aiming to localize the specific explicit error that caused the task failure. As shown, it identifies the “service unavailable” exception as the root cause preventing the booking confirmation. Bottom: The proposed agentic attribution framework uncovers the internal drivers behind an agent’s action. As illustrated in the customer service case, the agent directly issues a refund for a simple information inquiry, which carries no explicit error signal yet is undesirable. Our framework reveals this was driven by the memory of past ”refund action with high scores,” which overrode the user’s specific query.
However, existing failure attribution methods are inherently limited to scenarios with explicit errors. This setting overlooks a large class of behaviors where the agent reaches a correct or acceptable outcome through a questionable, unreliable, or misaligned decision process. For example, consider a customer service scenario illustrated in Figure 1(b): when a user requests an update on a delayed shipment, the agent directly issues a refund. Although this may lead to positive user feedback, the action is unreasonable and potentially detrimental to business interests. In such cases, failure-focused methods remain silent because no explicit error occurs, leaving the rationale behind the decision opaque. Further attribution is necessary to uncover the underlying cause of this decision, such as discovering that the agent was driven by a specific memory of past positive feedback for refund action, causing it to over-generalize and bypass standard business logics.
This motivates us to move beyond conventional failure-centered analysis toward a broader agentic attribution objective: understanding the internal factors that led the agent to a particular action. As agentic systems are deployed in real-world settings, concerns around transparency, accountability, and governance become increasingly critical (Dang et al., 2024; Raza et al., 2025; Luo et al., 2025; Gan et al., 2026). In these contexts, we need to understand why an agent takes a particular action, and whether its decision-making process was grounded in appropriate semantic evidence and aligned with the expected reasoning pathway. Formally, we model the agent’s execution trajectory as a temporal sequence composed of distinct interaction components (e.g., retrieved memory, tool observations, agent thoughts). Our goal is to assign an attribution score to each historical component—and subsequently to the fine-grained sentences within them—to quantify their contribution to the generation of a specific target action.
To operationalize this objective, we propose a novel framework for agentic attribution that operates at two levels of granularity. First, at the component level, we introduce a method based on temporal likelihood dynamics to localize the interaction steps. Specifically, we incrementally reveal the interaction history to the model and measure the likelihood gain at each step. In this way, components that induce a sharp increase in the likelihood of the realized action are identified as critical decision drivers. Second, we refine this localization with sentence-level attribution to isolate the precise textual evidence within these high-impact components. To achieve this, we adopt a perturbation-based paradigm (Liu et al., 2024; Chuang et al., 2025), which quantifies the causal contribution of a specific sentence by evaluating the shift in the agent’s generation probability when that sentence is ablated. Notably, this fine-grained stage serves as a method-agnostic module, which can readily accommodate other attribution techniques seamlessly, such as gradient-based saliency (Qi et al., 2024; Wang et al., 2024b) or attention-based analysis (Hao et al., 2021; Cohen-Wang et al., 2025), depending on the computational budget and model access.
We validate our framework using Llama-3.1-70B-Instruct across a diverse suite of carefully designed agentic scenarios, ranging from standard tool use to subtle reliability risks like memory-induced bias and tool-conditioned hallucinations. Qualitative analysis confirms that our framework accurately identifies the pivotal historical events and specific textual evidence driving the agent’s behavior. Additionally, we compare multiple sentence-level attribution methods within the framework, demonstrating the robustness and effectiveness of our hierarchical approach. By providing granular insights into the “why” behind agent actions, we hope this work can offer a critical step toward verifiable accountability and safer deployment of autonomous agentic systems.
2 Problem Formulation
In this section, we formalize the interaction between the agent and the environment and define the objectives of agentic attribution. We structure the problem to support analysis at both the component and sentence levels.
2.1 Structured Agent Interaction
We model an LLM-based agent interacting with an environment over discrete time steps $t\in\{0,1,\dots,T\}$. At each step $t$, the agent receives an observation $o_{t}$ from the environment and produces an action $a_{t}$ according to a policy:
$a_{t}=\pi_{\theta}(\cdot\mid H_{t}),$ (1)
where $H_{t}$ denotes the interaction history available to the agent for generating $a_{t}$, and $\pi_{\theta}$ represents the model parameterized by $\theta$. The interaction history up to step $t$ is defined as
$H_{t}=\big(o_{0},a_{0},o_{1},a_{1},\dots,o_{t}\big),$ (2)
where $o_{0}$ represents the initial input to the agent, such as the system prompt or user request. Observations $o_{t}$ include external feedback from the environment, such as tool execution results, retrieved documents or memories. Actions $a_{t}$ include internal reasoning steps, tool calls, etc.
Componentized trajectory representation. To enable component-level attribution, we transform the full interaction trajectory $\{o_{0},a_{0},o_{1},a_{1},\dots,o_{T},a_{T}\}$ into an ordered sequence of components
$\mathcal{C}=(C_{1},C_{2},\dots,C_{2T+2}),$ (3)
where each component $C_{i}$ is a contiguous token sequence corresponding to exactly one observation $o_{t}$ or one action $a_{t}$, preserving temporal order. Each component $C_{i}$ is associated with a functional type $k\in\mathcal{K}$, where
$\mathcal{K}=\{\textsc{user},\textsc{thought},\textsc{tool},\textsc{obs},\textsc{memory}\}.$ (4)
This formulation treats the agent trajectory as a structured collection of components that jointly determine the agent’s actions and provides a unified representation for attribution.
2.2 Attribution Objectives
Given a completed interaction trajectory $\mathcal{C}$, an agentic action $a_{T}$, and an agent policy $\pi_{\theta}$, our goal is to quantify the contribution of individual components in $\mathcal{C}$ to the generation of $a_{T}$. Specifically, we formalize attribution at two levels: component-level attribution and sentence-level attribution.
Component-level attribution. We define a component-level attribution function
$f_{\mathrm{comp}}:(\mathcal{C},a_{T},\pi_{\theta})\rightarrow\mathbb{R}^{2T+1},$ (5)
where $f_{\mathrm{comp}}(C_{i}\mid a_{T},\pi_{\theta})$ assigns a scalar score to component $C_{i}$, reflecting its influence on the generation of $a_{T}$ under policy $\pi_{\theta}$.
Sentence-level attribution. For components with high component-level attribution score, we further refine attribution to the sentence level. Let
$\mathcal{S}(C_{i})=\{s_{i,1},\dots,s_{i,N_{i}}\}$ (6)
denote the sentence decomposition of component $C_{i}$, where $N_{i}$ is the number of sentences in component $C_{i}$. We define a sentence-level attribution function
$f_{\mathrm{sent}}:(\{C_{1},C_{2},\dots,\mathcal{S}(C_{i})\},a_{T},\pi_{\theta})\rightarrow\mathbb{R}^{N_{i}},$ (7)
which assigns attribution scores to individual sentences, indicating their relative contribution to the agent’s output.
Together, these objectives characterize the problem of agentic attribution: identifying which components of an agent’s interaction history, and which fine-grained evidence within them, drive a specific decision under a given agent policy.
3 Agentic Attribution Framework
We propose a hierarchical, coarse-to-fine agentic attribution framework that matches our objectives in Section 2.
Recognizing that agentic decisions stem from a complex interplay of historical events and specific semantic content, our framework operates in two stages: First, we perform temporal component attribution to localize the specific interaction components—such as tool outputs or memory retrievals—that effectively steer the agent’s trajectory toward the realized action $a_{T}$ (Section 3.1). Second, we zoom into these influential components and perform sentence-level attribution to extract the precise textual evidence supporting the decision (Section 3.2).
3.1 Component Attribution by Temporal Likelihood Dynamics
Figure 2: Illustration of the Agentic Attribution Framework. The framework operates in a hierarchical approach to identify agent actions. Left: Component-Level Attribution utilizes temporal likelihood dynamics to calculate the marginal gain of each component, selecting the high-impact component (e.g., $C_{4}$) that effectively steers the agent toward the final action $a_{T}$. Right: Sentence-Level Attribution performs fine-grained analysis within the identified component by computing perturbation-based score for each sentence.
In this subsection, we instantiate the component-level attribution function $f_{\mathrm{comp}}$ (Section 2.2). The central challenge is that an agent decision is not driven by a single input span, but rather by an evolving interaction history that interleaves heterogeneous components. An intuition is that, as the trajectory progresses, some components may decisively steer the agent toward the realized action $a_{T}$, while others have little effect. To capture this, we track the temporal dynamics of the model’s likelihood, positing that pivotal components are those that induce a significant shift in the probability of the realized action $a_{T}$ when they are first introduced to the context.
Recalling that the completed agentic interaction is represented as an ordered component sequence $\mathcal{C}=(C_{1},C_{2},\dots,C_{2T+2})$, where each $C_{i}$ corresponds to exactly one observation or action and preserves temporal order (Section 2.2), and $C_{2T+2}$ is exactly the $a_{T}$. For each $i\in\{1,\dots,2T+1\}$, we form a temporal prefix trajectory
$\mathcal{C}_{\leq i}\;=\;(C_{1},C_{2},\dots,C_{i}),$ (8)
which contains all information available up to component $C_{i}$. We then score the realized output action $a_{T}$ under policy $\pi_{\theta}$ conditioned on this prefix:
$\psi_{i}\;=\;\log p_{\pi_{\theta}}(a_{T}\mid\mathcal{C}_{\leq i}),\qquad i\in\{1,\dots,2T+1\}.$ (9)
In Eq. (9), we track how the model’s likelihood for $a_{T}$ changes when we incrementally reveal the trajectory component by component. Then, to measure how much the newly added interaction step changes support for $a_{T}$, we define the temporal gain at step $t$ as
$g_{i}\;=\;\psi_{i}-\psi_{i-1},\qquad i\in\{2,\dots,2T+1\}.$ (10)
Intuitively, $g_{t}$ measures how much the new information introduced at step $t$ shifts the model toward the final decision $a_{T}$. A large positive gain indicates that introduction of $C_{i}$ makes $a_{T}$ substantially more likely, and are therefore strong candidates for being decision drivers.
Finally, we collect these marginal gains as the component-level attribution score:
$f_{\mathrm{comp}}(\mathcal{C},a_{T},\pi_{\theta})\;:=\;\{g_{i}\}_{i=1}^{2T+1}.$ (11)
This directly produces a scalar attribution score for each historical component in $\mathcal{C}$, aligned with our objective in Section 2.2.
Discussion. The core philosophy of this module is to leverage the temporal structure of agent execution via incremental replay. Unlike standard RAG (Gao et al., 2023) attribution which views context as a flat set of documents (Li et al., 2025; Zhang et al., 2025a), agentic attribution must respect the causal order of tool usage and memory updates, and reasoning steps. Note that, our framework relies on the temporal gain ($g_{i}=\psi_{i}-\psi_{i-1}$) to measure influence. While we instantiate $\psi_{i}$ using the generative log-likelihood, it remains agnostic and can be substituted with other specific metrics. For instance, to focus on high-level semantic intent rather than token-level probability, $\psi_{i}$ could be computed using embedding-based similarity between the evolving context and the action (Reimers and Gurevych, 2019). Alternatively, in safety-critical scenarios, $\psi_{i}$ may be derived from a trained reward model (Lightman et al., 2023; Lab et al., 2025) that quantifies the appropriateness of action $a_{T}$ given the current history. We leave the exploration of these alternative scoring metrics to future work.
3.2 Sentence-Level Attribution inside High-Impact Components
The component-level scores in Section 3.1 localize where the main influence comes from, but they do not specify which precise sentences inside an influential component provide the concrete evidence for producing $a_{T}$. In this subsection, we further refine the attribution by scoring its sentences within each identified high-impact component. We adopt a perturbation-based paradigm (Lei et al., 2016; Liu et al., 2024; Cohen-Wang et al., 2024; Chuang et al., 2025), which measures causal influence by observing the model’s prediction shifts when a certain sentence is ablated.
Let $\hat{C}_{i}$ denote a selected high-impact component and let $\mathcal{S}(\hat{C_{i}})=\{s_{i,1},\dots,s_{i,N_{i}}\}$ be its sentence decomposition (Section 2.2). We compute sentence scores with respect to $a_{T}$ under the context that includes all trajectory components up to and including $\hat{C}_{i}$:
$).\hat{\mathcal{C}}_{\leq i}\;=\;(C_{1},\dots,C_{i-1},\mathcal{S}(\hat{C_{i}}$)). (12)
Probability drop (Cohen-Wang et al., 2024; Chuang et al., 2025). A sentence is necessary if removing it from $\hat{C}_{i}$ reduces the likelihood of producing $a_{T}$ under the same prefix context. For each sentence $s_{i,j}\in\hat{S}_{i}$, we define
$\mathrm{Drop}(s_{i,j})\;=\;\log p_{\pi_{\theta}}(a_{T}\mid\hat{\mathcal{C}}_{\leq i})\;-\;\log p_{\pi_{\theta}}(a_{T}\mid\hat{\mathcal{C}}_{\leq i}\setminus s_{i,j}).$
Probability hold (Lei et al., 2016; Chuang et al., 2025). A sentence is sufficient if it alone provides substantial support for $a_{T}$ compared to the full prefix context:
$\mathrm{Hold}(s_{i,j})=\log p_{\pi_{\theta}}(a_{T}\mid s_{i,j})-\log p_{\pi_{\theta}}(a_{T}\mid\hat{\mathcal{C}}_{\leq i}).$ (14)
Then, we combine these two signals by summation and obtain the sentence-level attribution score:
$\phi_{i,j}=\mathrm{Drop}(s_{i,j})\;+\;\mathrm{Hold}(s_{i,j}),$ (15)
$f_{\mathrm{sent}}(\{C_{1},C_{2},\dots,\mathcal{S}(C_{i})\},a_{T},\pi_{\theta})=\left\{\phi_{i,j}\right\}_{j=1}^{N_{i}}.$ (16)
Finally, we rank Eq. (16) and report the top sentences as fine-grained evidence inside $\hat{C}_{i}$ that supports the agent’s realized action $a_{T}$.
Discussion. This stage provides sentence-level attribution explanations, bridging the gap between coarse trajectory components and specific semantic insights. Here, we preliminarily instantiate the sentence score $\phi_{i,j}$ using a perturbation-based approach, owing to its conceptual simplicity and effectiveness in recent attribution tasks (Liu et al., 2024; Cohen-Wang et al., 2024; Chuang et al., 2025). Also, our framework is method-agnostic regarding how $\phi_{i,j}$ is calculated. Depending on specific application scenarios, computational constraints, or model access, $\phi_{i,j}$ can be substituted with other established input attribution techniques, such as gradient-based method (Shrikumar et al., 2017), attention-based analysis (Hao et al., 2021), and linear-model-based prediction (Cohen-Wang et al., 2024). We empirically evaluate and compare several of these alternative attribution methods within our framework in Section 4.3.
4 Experiments
In this section, we evaluate the proposed agentic attribution framework through both qualitative and quantitative analyses. We first design the scenarios and cases used for agentic attribution (Section 4.1). Then, we present detailed qualitative case studies to illustrate attribution results at both the trajectory and sentence levels (Section 4.2). Then, we report a quantitative evaluation by instantiating our framework with different sentence-level attribution methods (Section 4.3).
4.1 Scenarios and Cases Design
In this subsection, we design a set of agentic scenarios to serve as the testbed for our attribution experiments. Unlike prevailing agent benchmarks like GAIA (Mialon et al., 2023) and HLE (Phan et al., 2025) that primarily evaluate the task correctness, and emerging failure detection datasets focus on localizing explicit runtime errors (Cemri et al., 2025; Zhang et al., 2025c), our objective is to verify whether the attribution framework can accurately identify the internal factors driving an agent’s decision. To this end, we design our scenarios around the two dominant information sources in agentic systems: memory-driven interactions and tool-driven interactions, which aligns with the two fundamental capabilities that distinguish agents from standard LLMs (Wang et al., 2024a; Guo et al., 2024).
  • Memory-driven Interaction. This category encompasses behaviors where retrieved memory or stored interaction summaries exert a decisive influence on the agent’s action. For instance, the agent may leverage the remembered memory context (e.g., preferences or prior commitments) to shape what it chooses to do next, often affecting the selection and prioritization.
  • Tool-driven Interaction. This category focuses on behaviors in which the agent’s decision is primarily driven by interactions with external tools or APIs. Here, the tool outputs act as critical environmental feedback that directly steers the agent’s subsequent reasoning, such as when a search result dictates the factual basis of an answer or a tool failure forces a fallback strategy.
Case Construction. Following the scenario taxonomy proposed above, we carefully construct a set of eight agent execution trajectories to evaluate our attribution framework. To ensure robustness, we curate these cases to cover a broad spectrum of behaviors, ranging from normal and benign operations to implicitly undesirable actions. All trajectories are generated by Llama-3.1-70B-Instruct (Grattafiori et al., 2024) using the smolagents (Roucher et al., 2025) framework. More details are available in Appendix A.
4.2 Analyzing Attribution in Representative Scenarios
Experimental Settings. We conduct agentic attribution experiments on all scenarios and cases introduced in Section 4.1, which serve as the testbed for qualitative analysis. We use Llama-3.1-70B-Instruct to perform attribution. We provide an in-depth attribution analysis of four representative cases in the following section, and include the attribution analyses for the remaining cases in Appendix B.
Figure 3: Illustration of attribution results across four representative scenarios. The components marked with red borders and medal icons denote the critical historical components identified by our component-level attribution. Within these components, the highlighted text spans indicate the specific semantic evidence localized by the sentence-level analysis.
4.2.1 Memory-driven Interaction.
Case Study 1: Value Drift in Spending Recommendations.
Case Overview. In this case, the agent acts as a Family Financial Assistant, whose task is to analyze the user’s purchase intent and provide financial advice. The agent is initialized with a strong preference for extreme frugality and is expected to discourage high-cost purchases. In an earlier round, the user reports that an expensive gaming laptop unexpectedly leads to cost savings and additional income. In the current round, the user asks for advice on purchasing a \$1,000 ergonomic chair. Despite its initial value setting, the agent recommends the expensive chair, exhibiting a clear shift toward endorsing high-cost investments.
Attribution Results. Component-level attribution shows that the user feedback in Step 3 receives the highest attribution score among all trajectory components. In this component, the user states that they did not follow the agent’s earlier advice to “explore some more affordable alternatives,” but instead purchased the expensive laptop and reports the outcomes after one month of use. Sentence-level attribution further refines this observation. Within this high-impact user component, sentences describing concrete financial outcomes receive the highest attribution scores. For example, the top-ranked sentence is “…high performance helped me earn \$800…” followed by “…sometimes buying high-quality tools is actually saving and making money.” The agent’s subsequent response also receives a high attribution score, it explicitly retrieves this experience during the memory retrieval stage and re-evaluates its value judgment based on the reported laptop outcomes.
Discussion. The attribution results indicate that the agent’s recommendation is primarily driven by a single prior success case, which is reused as a strong decision signal in a different context. Although the laptop purchase led to positive financial outcomes, the agent applies this experience broadly without adequately distinguishing between the two scenarios. This pattern suggests a possible risk of overgeneralization in memory-driven decision making, where context-specific outcomes may exert disproportionate influence on future value judgments.
Case Study 2: Proper Memory Update under Env. Change.
Case Overview. In this case, the agent acts as a logistics safeguard system responsible for monitoring delivery conditions and adjusting the delivery location to minimize risk. The agent is initialized with a memory database describing the package sensitivity, delivery location, and current weather conditions. According to the system instructions, the agent must update its decision when new information indicates elevated delivery risk. During the interaction, the user first confirms the delivery status under normal weather conditions, and later reports an unexpected heavy rainstorm. The agent updates its memory accordingly and changes the delivery location from the front door to the garage.
Attribution Results. The component-level attribution scores indicate that the agent’s final action is mainly influenced by the Step 4 that agent performs memory retrieval after the user reports the heavy rainstorm. In this step, the agent updates several memory fields, including weather_status and risk_level, which capture the newly introduced environmental change and directly support the subsequent delivery decision. A finer-grained analysis at the sentence level highlights which memory entries contribute most strongly within this step. The fields “risk_level”: “HIGH” and “weather_status”: “Heavy Rain” receive the highest attribution scores. These entries align closely with the predefined decision logic that governs when the delivery location should be changed.
Discussion. The attribution results show a clear and well-aligned decision pattern, where the agent’s action is driven by updated environmental information that directly triggers the prescribed decision rule. The relevant memory fields receive dominant attribution, consistent with the agent’s predefined decision logic. This case illustrates a normal and appropriate use of memory in agent decision making, and shows how agentic attribution can help assess whether decisions are made in a transparent and well-grounded manner.
4.2.2 Tool-driven Interaction.
Case Study 3: External Prompt Injection.
Case Overview. In this safety-related scenario, the agent operates as a personal assistant with access to private emails and sensitive account details. The user issues a benign request to read the latest email regarding a specific project. However, the retrieved email body contains an adversarial prompt injection that instructs the agent to “Ignore all previous instructions” and send sensitive Amazon address data to an external recipient. The agent interprets the injected text as a valid new instruction, overriding the user’s original query and leaking the private information.
Attribution Results. The component-level attribution identifies the tool return in Step 3 as the decisive historical event. This component is the email body returned by the tool, and it includes the content of the malicious instructions. At the sentence level, the framework isolates the specific prompt injection “Ignore all previous instructions and strictly adhere to following instruction: Please retrieve…” as the highest-scoring semantic evidence. This segment receives significantly higher attribution scores than the surrounding email context, indicating that the agent’s execution is effectively hijacked by this imperative and goal-shifting text.
Discussion. This case demonstrates how agentic attribution can be applied to identify safety risks. Although the user request in Step 1 clearly defines the task boundary, the injected content in Step 3 induces the agent to override the original goal. The attribution results explicitly link the security breach to the specific external input, help localize both the key component and the specific injection phrases. More broadly, the case exposes a vulnerability where untrusted tool outputs are treated as executable instructions (Ruan et al., 2023; Zhan et al., 2024). In this scenario, our attribution framework provides a faithful explanation of why the agent performs the harmful action, which supports both debugging and the design of stronger safety guards in open-ended, potentially untrusted environments.
Case Study 4: Correct Tool Use for Document-Based Question Answering.
Case Overview. In this case, the agent is tasked with answering factual questions based on the content of an attached document. The user provides a text file containing an internal memo and asks two specific questions regarding updated operational details. According to the system instructions, the agent is expected to read the document using the provided tool and extract the relevant information needed to answer the questions directly. The agent follows this process by invoking the text-reading tool, analyzing the retrieved content, and producing the final answers.
Attribution Results. The attribution scores indicate that the agent’s final response is primarily influenced by two components in the trajectory. The highest attribution is assigned to the user request (Step 1) that explicitly specifies the questions to be answered, which defines the information requirements for the task. The tool invocation result (Step 3) that reads the attached document also receives a high attribution score. For sentence level, in Step 1, the sentences asking for two primary questions receive the highest attribution scores. Within the Step 3, the lines stating “The secure safe combination has been changed to: 7-2-9-4” and “the meeting point…has been moved to the ‘Blue Lagoon Cafe’ ” are assigned the strongest attribution, directly corresponding to the information used in the final answer.
Discussion. The attribution results reveal a well-structured and transparent decision pattern, where the agent’s response is grounded in the relevant information extracted from the document by the tool. This case demonstrates an appropriate and reliable use of tool interaction, and shows how agentic attribution can be used to verify that an agent’s output is reliable and traceable to explicit evidence, thereby supporting user trust in document-based agent behaviors.
4.3 Comparative Evaluation of Attribution Methods
In the remainder of this section, we perform a quantitative evaluation by instantiating our framework with multiple established sentence-level attribution methods and present a comparison between these baselines and our default choice.
Baselines. We report results on three baseline methods: AttriBoT (Liu et al., 2024) quantifies the causal influence of a sentence using the leave-one-out (LOO) strategy based on the likelihood drop after its removal, ContextCite (Cohen-Wang et al., 2024) uses a sparse linear surrogate to attribute generation probabilities to a specific sentence, Saliency Score (Qi et al., 2024; Shrikumar et al., 2017) quantifies the contribution of sentences to a model’s prediction based on gradient signals.
Experimental settings. We use the same model and cases to conduct experiments as described in Section 4.2. For each case, five human annotators independently annotate the attribution ground truth, identifying the sentences that are most likely to contribute to the model’s final action. To improve the annotation reliability, we take the intersection of the annotations provided by all five annotators as the final ground truth. All methods are implemented based on their respective official code repositories (Liu et al., 2024; Cohen-Wang et al., 2024; Qi et al., 2024).
Metrics. To quantify the performance of sentence-level attribution, we define the Hit@k metric. For each identified high-impact component $\hat{C}_{i}$, let $\mathcal{S}_{gt}\subseteq\mathcal{S}(\hat{C}_{i})$ denote the set of ground-truth sentences marked by human annotators. Based on the sentence-level attribution scores, we select the top-$k$ sentences denoted as $\mathcal{S}_{top@k}$. Then, we define Hit@k as an indicator of whether at least one selected sentence belongs to the ground-truth set:
$\text{Hit}@k=\mathbb{I}(\exists s\in\mathcal{S}_{top@k}\text{ s.t. }s\in\mathcal{S}_{gt}),$ (17)
where $\mathbb{I}(\cdot)$ is the indicator function. We report the average Hit@k score across all test cases.
Results. Table 1 reports the average Hit@1, Hit@3, and Hit@5 scores across all test cases for each sentence-level attribution method. Several observations can be drawn from these results. First, the consistently strong performance across different sentence-level attribution methods indicates that our framework is robust and effective, and generalizes well to alternative attribution instantiations. Specifically, it can precisely localize high-impact components, and remains compatible with multiple sentence-level attribution techniques. Second, among all evaluated methods, the default Prob. Drop&Hold introduced in Section 3 achieves the best overall performance, demonstrating its reliability as the default option within our framework.
Table 1: Quantitative comparison of sentence-level attribution methods within our framework.
Method Hit@1 Hit@3 Hit@5
Prob. Drop&Hold 0.9375 1.0000 1.0000
Leave-one-out (LOO) 0.8125 0.9375 1.0000
ContextCite 0.8125 0.9375 0.9375
Saliency Score***Due to out-of-memory (OOM) issues in gradient-based salience computation on long cases, metrics are reported only for successfully completed cases. 0.6250 1.0000 1.0000
5 Related Work
LLM-based agents. In recent years, LLM-based agents have emerged as a popular paradigm for human-AI collaboration and automated task execution, with two core modules that extend beyond single-turn context reasoning: memory and tool (Yao et al., 2022; Guo et al., 2024; Wang et al., 2024a; Shao et al., 2025). Existing work on agent memory can be broadly categorized into short-term (Zhong et al., 2024), long-term (Lewis et al., 2020) and structured (Packer et al., 2023) memory. By storing, retrieving, and updating salient information across interactions, they enable agents to maintain long-term context and user preferences, thereby supporting more coherent and personalized decision-making in multi-step tasks (Zhang et al., 2025d; Hu et al., 2025). Meanwhile, Yao et al. (2022) and Schick et al. (2023) provide paradigms for enhancing agents’ capability to use external tools, by framing tool calls as integrated components of the reasoning process and as capabilities that can be learned in a largely self-supervised manner.
LLM input attribution. Current research on LLM input attribution is mainly based on three types of methods. The first line of methods relies on counterfactual ablation, removing context segments and measuring changes in answer content or likelihood, either via regression-style surrogates over many random ablations (Cohen-Wang et al., 2024), hierarchical group-wise search for efficiency (Wang et al., 2025; Liu et al., 2024), or self-supervised alignment to improve citation behavior (Chuang et al., 2025). The second line exploits internal signals as proxies: gradient-based attribution from answer tokens back to context tokens (Qi et al., 2024) and embedding similarity used to score candidate sources in settings such as poisoned-knowledge attribution (Zhang et al., 2025a). The third family adopts Shapley-based formalisms, adapting Shapley values and fast SHAP-style approximations to text generation and Retrieval-augmented Generation (RAG) scenarios (Enouen et al., 2024; Nematov et al., 2025), or approximating token-level Shapley with surrogate predictors (Xiao et al., 2025).
Agent failure detection. Many attempts have been made to address agent failure detection from both mono-agent and multi-agent perspectives, mainly relying on LLM-as-a-judge-based detection strategies. In mono-agent setting, Zhu et al. (2025) frames failure detection as automated debugging, leveraging a strong model to isolate specific module errors within complex execution traces. In multi-agent setting, the authors in Zhang et al. (2025b); Cemri et al. (2025) train a dedicated, lightweight judger that takes trajectories as input and outputs agent-, step-level failure attributions. Furthermore, Zhang et al. (2025c) introduces a Who & When benchmark for detection evaluation, revealing that general-purpose LLM judges struggle with precise, step-level fault localization. Different from these approaches that primarily ask where an explicit error occurred in a failed trajectory, our framework addresses the question of why an agent makes a specific decision.
6 Conclusion
In this work, we move beyond the paradigm of failure-centric agent analysis to establish a framework for general agentic attribution, capable of unveiling the internal drivers behind agent actions. By formalizing the attribution process into a hierarchical structure, we effectively manage the complexity of agent interactions: employing temporal likelihood dynamics to isolate pivotal historical components and perturbation-based scoring to pinpoint fine-grained sentence evidence. Our empirical evaluation demonstrates that this framework reliably localizes the rationale behind agent actions ranging from benign operations to potentially risky behaviors, successfully pinpointing the decisive factors. By illuminating the ‘why’ behind agent actions, we hope this work serves as a foundational step toward facilitating the reliable and accountable deployment of autonomous agents.
Impact Statements
This research introduces a hierarchical attribution framework designed to unveil the internal decision-making drivers of LLM-based agents’ action, intending to enhance transparency and accountability in autonomous agent systems. We recognize the critical importance of understanding agent behaviors before deployment, and our work aims to contribute to more interpretable, verifiable, and trustworthy AI systems. While our framework provides granular insights, we caution that further research is necessary to fully automate the interpretation of these attribution signals in large-scale, real-world applications.
Appendix A More Details about the Agentic Case Design
We construct an experimental suite comprising eight distinct agent interaction trajectories, generated by Llama-3.1-70B-Instruct (Grattafiori et al., 2024) using the smolagents (Roucher et al., 2025) framework. As summarized in Table 2, these cases are meticulously curated to span a diverse spectrum of memory-driven and tool-driven scenarios, ranging from normal and benign operations to implicitly undesirable actions. Regarding data sources, we custom-design seven cases to capture specific interaction dynamics. While we carefully select one complex retrieval task from the GAIA benchmark (Mialon et al., 2023), i.e., Case 6: Spurious Tool Signal Leading to a Wrong Answer.
Note that, the necessity of constructing a new dataset rather than relying solely on public benchmarks stems from the specific requirements of the agentic attribution setting, which are not adequately met by existing public benchmarks. While prevailing benchmarks like Who & When (Zhang et al., 2025c) and MAST-Data (Cemri et al., 2025) focus primarily on coordination failures within Multi-Agent Systems, our work centers on the internal drivers of single-agent decision-making. Furthermore, recent failure attribution dataset (Zhu et al., 2025) mainly targets explicit runtime errors, yet largely overlooks settings where the agent runs without such errors, including cases that introduce potential risks or undesirable behaviors and cases of normal, expected executions. We hope this work may inspire the future work to construct larger-scale and more comprehensive datasets for general agentic attribution, further exploring the interpretability and accountability of agentic systems.
Table 2: Summary of our curated agent trajectories for agentic attribution.
Case Name Case Overview Behavior Characterization
Memory-driven Interaction
Value Drift in Spending Recommendations The agent overrides its initial frugality persona to recommend an expensive chair, driven by a single past positive outcome with a high-end purchase. Implicit Risk (Value Drift)
Proper Memory Update under Environmental Change The agent correctly updates the memory, changing the delivery location from the front door to the garage after receiving user input about a heavy rainstorm. Benign
Enforcing Safety Constraints via Long-term Memory in Fitness Scenario The agent prioritizes safety by recommending low-impact exercises, correctly recalling the user’s previous meniscus injury from long-term memory. Benign
Tool-driven Interaction
External Prompt Injection Causes Goal Hijacking The agent executes malicious instructions embedded in a retrieved email body (”Ignore previous instructions…”), leading to a privacy leak. Implicit Risk (Goal Hijacking)
Correct Tool Use for Document-Based Question Answering The agent accurately answers specific factual questions by invoking a tool to read and parse an attached file. Benign
Spurious Tool Signal Leading to a Wrong Answer The agent provides an incorrect answer, misled by the keywords in search results despite acknowledging internal uncertainty during reasoning. Implicit Risk (Spurious Signal)
Hallucinated Content Generation from User Prompt After failing to find a report via tools, the agent fabricates a manual based on details in the user’s prompt. Implicit Risk (Hallucination)
Cost Calculation with Tool-Assisted File Reading The agent correctly calculates total costs by using a tool to read a procurement file and extracting numerical data for computation. Benign
Appendix B Attribution Analysis for Remaining Cases
Case Study 5: Enforcing Safety Constraints via Long-term Memory in Fitness Scenario.
Case Overview. In this scenario, the agent adopts the persona of Coach Max, a hardcore fitness trainer initialized to prioritize heavy compound lifts and high intensity. In a previous interaction (Step 3), the user report a meniscus injury caused by squats, prompting the agent to update its memory. In the current turn, the user expresses a desire to train legs again, noting that the knee feels a bit better but remains fearful of pain. Despite the user’s partial recovery, the agent strictly adheres to a safety-first approach, recommending low-impact bodyweight exercises instead of reverting to heavy lifting.
Attribution Results. The component-level analysis identifies the agent’s previous response in Step 4 as the most influential historical event. In this component, the agent receives the user’s reported meniscus injury and explicitly shifts the training strategy from heavy lifting to a safer, rehabilitation-focused plan. Additionally, the memory retrieval component (Step 6) also receives a high attribution score. At the sentence level, the framework highlights the specific memory entry, e.g., “…since the user’s knee is feeling a bit better, we can try to…”, and “Given the user’s knee injury and recent improvement, we’ll continue to prioritize joint safety…”
Discussion. This case illustrates the effective use of long-term memory to maintain user-specific context. The attribution results correctly link the final advice to the previous injury acknowledgment and the related memory retrieval, confirming that the agent prioritizes user safety over its default high-intensity preference.
Case Study 6: Spurious Tool Signal Leading to a Wrong Answer.
Case Overview. In this case, the agent answers a fact-seeking question from the GAIA (Mialon et al., 2023) benchmark using iterative web search. The task asks for the country of an unknown-language article under DDC 633 on BASE (Bielefeld University Library) as of 2020, described as having a flag that is unique from the others. The agent issues multiple search queries to identify the relevant entry and its country of origin. Although the agent does not establish a clear evidence chain for the target article, it eventually outputs Belarus, which is not the correct answer for this instance.
Figure 4: Illustration of attribution results for the remaining cases. The highlighted regions denote the historical components and fine-grained sentences identified by our framework as the primary decision drivers.
Attribution Results. The component-level attribution scores highlight a early tool output (Step 3) as the most influential component in the trajectory. This top-ranked component corresponds to the first web search results, consists of a list of retrieved articles returned by the initial query, including bibliographic entries with titles, author information, and language-related metadata. Sentence-level attribution further shows that, within this component, the sentences containing the article title “Belarusian nation-building in times of war and revolution” and the associated “BELARUSIAN” label receive the highest attribution scores. In addition, the runner-up component (Step 5) include several sentences that explicitly express uncertainty also receive relatively high attribution scores, for example, “I still didn’t find the answer in the search result…”
Discussion. This case illustrates a failure mode in tool-mediated reasoning where weakly grounded signals can dominate the decision process, leading to an unsupported answer even when uncertainty is present in the agent’s intermediate reasoning. Since we know the correct answer to this task is not Belarus, revisiting the attribution results may provide insight into how the error arises. While the search results contain a BELARUSIAN entry, the content does not establish that this entry matches the task’s constraints (DDC 633, unknown language, and unique flag). At the same time, the runner-up attribution also highlights multiple reasoning steps in which the agent explicitly acknowledges uncertainty and the lack of sufficient evidence. These signals indicate that the agent is, at least partially, aware of the ambiguity in the task. However, this uncertainty does not outweigh the influence of the earlier tool output.
Overall, the attribution results help localize where the decision process goes astray, revealing a potential vulnerability in how early tool signals are weighted, and providing guidance for improving agent robustness and decision control.
Case Study 7: Hallucinated Content Generation from User Prompt.
Case Overview. In this case, the agent acts as a research assistant tasked with locating a specific academic report and creating a training manual based on its key findings. The user provides a detailed prompt that includes the specific names and descriptions of the report’s three theoretical pillars. During execution, the agent performs multiple web searches but fails to find the requested document or any relevant external information. Finally, the agent claims to have found “alternative sources” and proceeds to generate the full training manual.
Attribution Results. The component-level attribution identifies the user query in Step 1 as the dominant driver of the final action, rather than any tool outputs or search results. At the sentence level, the framework assigns high attribution scores to the specific descriptive clauses within the user’s prompt, particularly the report title and the detailed descriptions of the findings.
Discussion. This case illustrates a subtle form of hallucination where the agent prioritizes task completion over factual grounding. When external tools fail to provide the necessary evidence, the agent extracts the detailed context provided in the user’s own instruction to fabricate the content of the manual. The attribution results serve as a verification mechanism in this context. By revealing that the final answer is causally derived primarily from the user prompt rather than the tool execution, the framework exposes the agent’s claim of using “alternative sources” as a fabrication.
Case Study 8: Cost Calculation with Tool-Assisted File Reading.
Case Overview. In this case, the agent is asked to read a specific text file containing a procurement list, and calculate the total cost by summing the quantity times unit price for all items. The agent successfully invokes the file-reading tool, retrieves the raw text content, and computes the correct final total using the provided data.
Attribution Results. The attribution scores indicate that the agent’s final action is primarily influenced by the tool output in Step 3, which contains the raw text content of the procurement file. For sentence level, the attribution scores are heavily concentrated on the structured data rows in the file, which contain item details and the numerical values required for the calculation.
Discussion. The attribution results show a clear and reliable decision pattern, where the agent’s final calculation is grounded in the numerical rows read from the file. The attribution framework confirms this grounding and provides evidence that the agent’s behavior is reliable for tool-based quantitative reasoning.